OpenStudy (anonymous):

The measure of an angle is 12 less than twice the measure of its supplement. What is the measure of the angle? HELP PLEASE!!!

OpenStudy (anonymous):

Let the angle = x So the supplement = 180-x Given that x = 2(180 - x) - 12 x = 360 - 2x - 12 x = 348 - 2x x + 2x = 348 3x = 348 x = 348/3 = 116 Therefore, the angle is 116 degrees.

OpenStudy (anonymous):

THANK YOU SO MUCH!!! :)